OpenStudy (anonymous):

Help me study for my math final exam. Please. If I get less than 90% my parents will freak :c... the units we're covering is integers, ratios, percents, probability, pythogorean relation, linear relations, surface area and volume (3-D geometry) ^^ HELPPP ~ thx.

mathslover (mathslover):

what is the problem you are having any question ?

mathslover (mathslover):

be confident dear u will get above this 90% ....... i always get tensed but i always got 95% above :) ... now i understood overtension is not good ..but tension is necessary ;)

OpenStudy (anonymous):

Proportions. And unit rates. Help me idk if I can give example.

mathslover (mathslover):

see if u have to check whether a fraction can be further simplified or not like ... 102/200 = 56/100 = 28/50 = 14/25 try to find the HCF ( highest common factor of 14 and 25) GCF of 14 and 24 is 1 hence it can not be further simplified

mathslover (mathslover):

in ratios u can write any ratio as in a fraction form .. for example 2:3 = 2/3 111:1 = 111/1 = 111 proportions mean x:4::1:2 this means that 4*1 = 2*x 4=2x x=4/2 = 2 hence x= 2

mathslover (mathslover):

any other questions ?

OpenStudy (anonymous):

Aaaah okay I see. Also percents. Students have different ways to go to school: 100 students walk, 40 students get driven, 50 driven by bus and 20 take bike or other. What percent of students are driven. What percent walks? \[100+40+50+20 = 210\] (total) percent formula for this question = ?

mathslover (mathslover):

total no. of students = 210 right ?

mathslover (mathslover):

now : percentage of students who walk : students who walk / total students * 100 40/210 * 100% = 400/21= 19.5(approx.)%

OpenStudy (anonymous):

Oh okay. I get it. Next: Probability. How do you calculate theoritical and experimental? Other: Get a spinner with 4 colors reg green blue yellow. Toss a coin with every spin. What are the favourable and possible outcomes? What is the theoritical probab.?

mathslover (mathslover):

see : let me give u a simple example : we have a coin .. we tossed it what is the probabilty to get heads and tails total no. of observations = 2 heads :1 therefore probability of getting heads is 1/2 similarly for that one of tails..1/2

OpenStudy (anonymous):

and if you have a spinner with 4 colors?